I am having problems with this, can someone please help me with this question?
The length of a rectangular sign is 3feet longer than the width. If the sign has a area of 40 square feet for advertising, find its length and its width.

Answer by alka(15)

let width = x
length = x+3
area = length * width
40 = X(X+3)
40 = x^2+3x
x^2+3x-40=0
x^2+8x-5x-40=0
x(x+8)-5(x+8)=0
(x+8)(x-5)=0
x=-8 OR x=5
width cannot be negative
width=5feet
length=8feet